this is what i have could anyone help please
An insurance company has determined that in a certain region the probability of lightning striking a house in a given year is about 0.0005, and the average cost of repairs of lightning damage is $8000 per incident. The company charges $28 per year for lightning insurance.

A. What is the company's expected value for the net income from each lightning insurance policy in one year?
B. If the company has 300,000 lightning damage policies, what is the company's expected yearly income from lightning insurance?

Thanks guys

Found 2 solutions by jorel1380, Boreal:
Answer by jorel1380(3719)

A.The cost of repairs is $8000 per incident, while the probability of an incident is 0.0005. So the expected cost is 8000x0.0005, or $4 per policy. Thus, the net profit from one policy is 28-4, or $24/policy.
B. 300000 policies x $24/policy=$7,200,000 profit annually from lightning insurance. ☺☺☺☺

Answer by Boreal(15235)

E(x)=+28(.9995)-8000(.0005)=+27.986-4=+$23.986 or +$23.99
300,000*23.986=$7,195,800
======================
Check
300,000*.0005=150 with lightning damage and costs of $1,200,000
299,850*28=$8,395,800
The difference is $7,195,800